# Idiosyncrasies in Large Language Models

### v1

@@ B.4 Unique Markdown Formatting
B.5 Open-ended Language Analysis
@@ Appendix C Response Demonstrations
In this part, we present examples of LLM responses. Table 12 and 13 illustrate responses before and after our prompt-level interventions (Section 2.2). Table 14 shows the transformed responses when only markdown elements are retained (Section 3.2). Table 15 presents the rewritten responses (Section 3.3). Additionally, Table 16 provides a pairwise comparison of responses, supporting our findings in Table 9. Table 17,  18,  19,  20,  21,  22,  23,  24,  25,  26,  27,  28,  29 contain example responses from each LLM considered in Section 2, including ChatGPT, Claude, Grok, Gemini, DeepSeek, LLama3.1-8b (instruct), Gemma2-9b (instruct), Qwen2.5-7b (instruct), Mistral-v3-7b (instruct), LLama3.1-8b (base), Gemma2-9b (base), Qwen2.5-7b (base), and Mistral-v3-7b (base).

### v2

@@ 6 Conclusion
We conclude by outlining several directions for future work:
  • It is worth investigating whether our observations generalize to LLMs beyond the Transformer architecture, such as state space models (Gu & Dao, 2024) and diffusion-based language models (Nie et al., 2025).
  • Understanding how the training process results in these idiosyncrasies remains an important open question.
  • Third, our setup does not consider scenarios where the list of source LLMs could be large and even unknown beforehand, which require further investigation. Our initial results suggest it is promising: a 10-way classification problems involving 5 chat APIs and 5 instruct LLMs including Phi-4 achieves 92.2% accuracy.
  • It is interesting to examine how these idiosyncrasies relate to model distillation (Hinton et al., 2015), a technique that has become increasingly prevalent in practice.
@@ Acknowledgments
Impact Statement
Our study investigates the distinguishability of LLMs. On the positive side, our results offer insights into LLM behaviors that enhance transparency, accountability and provenance in tracking AI generated content. This has valuable applications such as combating misinformation, and ensuring responsible use of LLMs. However, our results highlight challenges in model training. Reliance on synthetic data or model distillation risks propagating biases and potentially infringing on the intellectual property of proprietary models.
@@ B.4 Unique Markdown Formatting
B.5 Rewriting LLM outputs
In Section 3.3, we used GPT-4o-mini to rewrite LLM outputs. Here, we present results using an alternative model: Qwen2.5-7B-Instruct. As shown in Table 12, our observations remain consistent, indicating that our findings are robust to the choice of LLM used for rewriting.
LLM for rewriting original paraphrase translate summarize
GPT-4o-mini 97.8 93.6 93.9 63.7
Qwen2.5-7B-Instruct 97.8 92.6 94.3 71.5
Table 12: Classification accuracies on rewritten responses. The results are on Chat API responses.
B.6 Open-ended Language Analysis
@@ Appendix C Response Demonstrations
In this part of the appendix, we present examples of LLM responses. Table 13 and 14 illustrate responses before and after our prompt-level interventions (Section 2.2). Table 15 shows the transformed responses when only markdown elements are retained (Section 3.2).
Table 16 presents the LLM rewritten responses (Section 3.3). For paraphrasing, we find that formatting style remains largely unchanged, e.g., the number of enumerated lists are the same. Most of the differences lie in their word choices, paraphrased texts use different words with similar meanings but do not change the high-level semantic meaning of the original texts.
Additionally, Table 17 provides a pairwise comparison of responses, supporting our findings in Table 9. Table 18,  19,  20,  21,  22,  23,  24,  25,  26,  27,  28,  29,  30 contain example responses from each LLM considered in Section 2, including ChatGPT, Claude, Grok, Gemini, DeepSeek, Llama3.1-8b (instruct), Gemma2-9b (instruct), Qwen2.5-7b (instruct), Mistral-v3-7b (instruct), Llama3.1-8b (base), Gemma2-9b (base), Qwen2.5-7b (base), and Mistral-v3-7b (base).